OpenStudy (anonymous):

find the constant of proportionality and the unit rate for the data in the table then write an equation to represent the relation ship between time t an distance d

OpenStudy (dayakar):

where is table

OpenStudy (michele_laino):

I think that the proportionality constant and the unit rate are the same quantity. In order to get such unit rate, please note that, if we compute the ratio between the distance and the corresponding time, we get \[\Large \frac{{270}}{6} = \frac{{225}}{5} = \frac{{135}}{3} = \frac{{90}}{2} = ...?\] please what is the value of such common ratio?

OpenStudy (anonymous):

0/1?

OpenStudy (michele_laino):

hint: what is \(90/2=...?\)

OpenStudy (michele_laino):

is \(90/2=45\)?

OpenStudy (anonymous):

so what is answer

OpenStudy (michele_laino):

please note that I can not give direct answers, since it is against the Code of Conduct

OpenStudy (anonymous):

so i would be 45

OpenStudy (michele_laino):

yes! that's right!

OpenStudy (michele_laino):

we have: \(45 \cdot 3=135\), \(45 \cdot 5=225\), and \(45 \cdot 6=270\)

OpenStudy (michele_laino):

therefore for a generic distance \(d\) and for a generic time \(t\), we can write: \[\huge \frac{d}{t} = 45\] Now, please if I multiply both sides of such equation by \(t\), I get: \[\huge \frac{d}{t} \cdot t = 45 \cdot t\] please simplify such equation, what do you get?

OpenStudy (michele_laino):

hint: 

OpenStudy (anonymous):

45

OpenStudy (michele_laino):

we have: \[\huge \frac{d}{t} \cdot t = d\] am I right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

well idk

OpenStudy (michele_laino):

so, after such simplification, we can write: \[\huge d = 45 \cdot t\] which is the requested relation between distance and time

OpenStudy (michele_laino):

and we have solved your exercise! :)